### v1

@@ D.2 Interactive Segmentation
Further, Fig. 10 shows detailed quantitative results for all the above mentioned datasets. We provide quantitative results for all trained PathoSAM generalist models. Overall, we observe that our generalist model clearly outperforms SAM and $\mu$SAM, and the segmentation quality is consistent along different variants of the generalist model. For some datasets, the model trained on PanNuke [Gamper et al.(2020)Gamper, Koohbanani, Benes, Graham, Jahanifar, Khurram, Azam, Hewitt, and Rajpoot] performs a bit better for iterative refinement, a likely reason could be the models’ distribution proximity to the training data.
@@ D.3 User Study
Here, we choose a H&E stained histopathology tissue image for validating SAM and PathoSAM for user-based interactive segmentation in two different tools. QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] expects the installation of SAM-API, which is well documented at https://github.com/ksugar/samapi. After starting the SAM-API server, the user can access the SAM extension and begin annotating. The tool currently supports default SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and SAM2 [Ravi et al.(2024)Ravi, Gabeur, Hu, Hu, Ryali, Ma, Khedr, Rädle, Rolland, Gustafson, Mintun, Pan, Alwala, Carion, Wu, Girshick, Dollár, and Feichtenhofer] models. If desired, users can provide the URL for remotely stored model checkpoints, which makes it compatible with SAM-family models. Fig. 11 displays a screenshot on top for annotation with default SAM model, and bottom for annotation with PathoSAM model. We clearly see that the model identifies nuclei better with our PathoSAM generalist model, using a single point or box per object. $\mu$SAM comes with the installation package for PathoSAM, also well documented at https://computational-cell-analytics.github.io/micro-sam/ for stand-alone usage. A user has the flexibility to start napari [Sofroniew et al.(2024)Sofroniew, Lambert, Bokota, Nunez-Iglesias, Sobolewski, Sweet, Gaifas, Evans, Burt, Doncila Pop, Yamauchi, Weber Mendonça, Buckley, Vierdag, Royer, Can Solak, Harrington, Ahlers, Althviz Moré, Amsalem, Anderson, Annex, Boone, Bragantini, Bussonnier, Caporal, Eglinger, Eisenbarth, Freeman, Gohlke, Gunalan, Har-Gil, Harfouche, Hilsenstein, Hutchings, Lauer, Lichtner, Liu, Liu, Lowe, Marconato, Martin, McGovern, Migas, Miller, Muñoz, Müller, Nauroth-Kreß, Palecek, Pape, Perlman, Pevey, Peña-Castellanos, Pierré, Pinto, Rodríguez-Guerra, Ross, Russell, Ryan, Selzer, Smith, Smith, Sofiiuk, Soltwedel, Stansby, Vanaret, Wadhwa, Weigert, Windhager, Winston, and Zhao] and select the Segment Anything for Microscopy available under extensions and a desired annotator (for histopathology images, the relevant choice is ”Annotator 2D”), or access the 2D annotator with CLI scripts. For images larger than the training data, which is quite common for WSIs, $\mu$SAM supports segmentation over tiles over the image. We make use of this feature to evaluate both SAM and PathoSAM models. In addition, we provide scripts to run segmentation over WSIs.

### v2

@@ D.2 Interactive Segmentation
Further, Fig. 14 shows detailed quantitative results for all the above mentioned datasets. We provide quantitative results for all trained PathoSAM generalist models. Overall, we observe that our generalist model clearly outperforms SAM and $\mu$SAM, and the segmentation quality is consistent along different variants of the generalist model. For some datasets, the model trained on PanNuke [Gamper et al.(2019)Gamper, Koohbanani, Benes, Khuram, and Rajpoot] performs a bit better for iterative refinement, a likely reason could be the models’ distribution proximity to the training data.
@@ D.3 User Study
Here, we choose a H&E stained histopathology tissue image for validating SAM and PathoSAM for user-based interactive segmentation in two different tools. QuPath [Bankhead et al.(2017)Bankhead, Loughrey, Fernández, Dombrowski, McArt, Dunne, McQuaid, Gray, Murray, Coleman, James, Salto-Tellez, and Hamilton] expects the installation of SAM-API, which is well documented at https://github.com/ksugar/samapi. After starting the SAM-API server, the user can access the SAM extension and begin annotating. The tool currently supports default SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, Dollar, and Girshick] and SAM2 [Ravi et al.(2025)Ravi, Gabeur, Hu, Hu, Ryali, Ma, Khedr, Rädle, Rolland, Gustafson, Mintun, Pan, Alwala, Carion, Wu, Girshick, Dollar, and Feichtenhofer] models. If desired, users can provide the URL for remotely stored model checkpoints, which makes it compatible with SAM-family models. Fig. 15 displays a screenshot on top for annotation with default SAM model, and bottom for annotation with PathoSAM model. We clearly see that the model identifies nuclei better with our PathoSAM generalist model, using a single point or box per object. $\mu$SAM comes with the installation package for PathoSAM, also well documented at https://computational-cell-analytics.github.io/micro-sam/ for stand-alone usage. A user has the flexibility to start napari [Sofroniew et al.(2024)Sofroniew, Lambert, Bokota, Nunez-Iglesias, Sobolewski, Sweet, Gaifas, Evans, Burt, Doncila Pop, Yamauchi, Weber Mendonça, Buckley, Vierdag, Royer, Can Solak, Harrington, Ahlers, Althviz Moré, Amsalem, Anderson, Annex, Boone, Bragantini, Bussonnier, Caporal, Eglinger, Eisenbarth, Freeman, Gohlke, Gunalan, Har-Gil, Harfouche, Hilsenstein, Hutchings, Lauer, Lichtner, Liu, Liu, Lowe, Marconato, Martin, McGovern, Migas, Miller, Muñoz, Müller, Nauroth-Kreß, Palecek, Pape, Perlman, Pevey, Peña-Castellanos, Pierré, Pinto, Rodríguez-Guerra, Ross, Russell, Ryan, Selzer, Smith, Smith, Sofiiuk, Soltwedel, Stansby, Vanaret, Wadhwa, Weigert, Windhager, Winston, and Zhao] and select the Segment Anything for Microscopy available under extensions and a desired annotator (for histopathology images, the relevant choice is ”Annotator 2D”), or access the 2D annotator with CLI scripts. For images larger than the training data, which is quite common for WSIs, $\mu$SAM supports segmentation over tiles over the entire image. We make use of this feature to evaluate both SAM and PathoSAM models. In addition, PathoSAM provides a simple CLI script to run automatic segmentation over WSIs.